FY1303 Elektricity and magnetism Spring + fall 2004

• Credit: 15 (“studiepoeng”)

• Duration: Two semesters.

• Lectures:

– Spring: 3 hours pr week

– Fall: 2 hours pr week

• Assignments (homework + classroom guidance):

– Spring: 2 hours pr week

– Fall: 1 hour pr week

• Laboratory: 7 exercises (ca 70 hours, including two reports)

• Compulsory activities:

– All laboratory exercises must be completed.

– At least 2/3 of the homework assignments must be completed.

• Evaluation forms:

– Midterm exam (March): 20%

– Midcourse exam (May): 40%

– Project work (incl report) (October/November): 20%

– Final exam (December): 20%

• Suggested textbooks: Alonso and Finn Physics, or similar (e.g. Young and Freedman; Tipler and Mosca; Fishbane, Gasiorowicz, and Thornton; Serway and Jewett; Knight)

Contents:

This is an introductory course in electromagnetism and contains: Electrostatics (Coulomb’s law, Electric fields and forces, Gauss’ law, Electric potential and energy, Conductors, Dielectrics, Capacitance); Magnetostatics (Magnetic fields and forces, Magnetic dipoles, Biot-Savart law, Ampere’s law, Magnetic flux, Magnetic materials); Electromagnetic induction (Faraday’s law, Lenz’ law, Inductance); Electric circuits (DC- and AC- circuits, Impedance, Resonance). The second half of the second semester is devoted to a project where the student selects a topic related to electromagnetism and works independently or in a small group. The project may be experimental or theoretical in character and involves writing a report.

Jon Andreas Støvneng Ass. professor Deparment of physics, NTNU